I made a mistake with the original question, here is the correction;
The formula G=~(2.5x+357) and L=7.43~(x-115) model the loan amounts, in billions of dollars, to college students for grants, G, and loans, L, x years after 1970. In which year was the amount from loans the same as the amount from grants? Please show your work.

Answer by stanbon(48564)

The formula G=~(2.5x+357) and L=7.43~(x-115) model the loan amounts, in billions of dollars, to college students for grants, G, and loans, L, x years after 1970. In which year was the amount from loans the same as the amount from grants? Please show your work.
----------
Loans will equal grants when G = L:
---------
EQUATION:
2.5x+357=7.43-(x-115)
2.5x+357=7.43-x+115
3.5x=7.43+115-357
3.5x=-234.57
x=-67.02
1970-67.02=1902
---------
I doubt that this is the correct answer but
it is the answer for what you have posted.
--------
Check your problem again; did you post it correctly?
Cheers
Stan H.